# An Survey of the Evolution of Language Model-Based Dialogue Systems

### v1

@@ 4.3 Fusion between LLM and DM
In summary, as depicted in Table IV, there is a discernible temporal trend towards an increased prevalence of open-sourced LLMs, particularly within the domain of dialogue models. This trend is accompanied by a continuous expansion in the scale of both training data and parameter configurations. The scaling law still works for current LLMs but most of released models is 7B or 13B, due to the computing cost of inferences. Furthermore, there is a noteworthy extension in the window length, denoting an enhanced ability to process and generate longer sequences of text. In the pursuit of improving the utility and safety of dialogue models, there is an evident integration of a higher volume of instruction fine-tuning using conversational data. Additionally, there is a consistent evolution in the multilingual capabilities exhibited by these models.
5 LLM-based Dialogue System
At present, passing the Turing Test and even simulating believable human behavior [192], such as waking up, cooking breakfast, and heading to work, is not a formidable challenge for LLM-based dialogue systems [193, 152]. However, this progress introduces new problems and challenges. First of all, due to unaffordable computing cost during training of LLMs, prompting has emerged as a predominant method for eliciting the reasoning capabilities of LLMs across various tasks [194, 195], including dialogue scenarios [5, 196] (§ 5.1). Secondly, the dialogue system not only interacts with the user but also interacts with external knowledge sources or tools, acquiring necessary intermediate results to resemble the final responses (§ 5.2). These two critical capabilities have become increasingly attractive and crucial for empowering dialogue systems, which can be utilized in complementary to suit the diverse and complicated needs in practice (§ 5.3). Table 4 shows the three major conversational flows for the LLM-based dialogue system.
@@ 5 LLM-based Dialogue System
Figure 4: A novel unified perspective of LLM-based Dialogue System. There are three conversational flows: 1) response based on internal reasoning and dialogue context; 2) response based on external sources and dialogue context; and 3) response based on internal reasoning, external sources, and dialogue context.
5.1 Internal Reasoning
@@
TABLE V: Summary of research works exploring LLMs for various dialogue tasks.
Method(s) Cue(s)† Type Task(s)* LLM(s) Settings Language Domain
Pan et al. [197] S TOD NLU ChatGPT ZS EN General
Michael et al. [198] S DST ChatGPT ZS EN General
IT-LLM [199] S DST, NLG ChatGPT ZS, FS EN General
SGP-TOD [187] S DST, DPL, NLG ChatGPT ZS EN General
GDP-ZERO [200] S DPL ChatGPT FS EN PersuasionForGood
Qin et al. [201] S ODD DS ChatGPT ZS EN General
ProCoT [196] S PD ChatGPT, Vicuna ZS, FS EN General
Fan et al. [202] S DDA ChatGPT ZS, FS CN, EN General
Huynh et al. [203] S DE GPT3, BLOOM FS EN General
Mendonça et al. [204] S DE ChatGPT ZS Multilingual General
Bang et al. [205] S TOD, ODD DRG ChatGPT ZS EN General
CueCoT [5] Per, Psy, E ODD DQA, DRG ChatGPT, ChatGLM ZS, FS CN, EN General
Zhao et al. [206] E DU, DRG ChatGPT ZS EN General
SMILE [207] Psy ODD DRG CN Mental Health
ExTES [208] E ODD DRG LLaMA Finetuning EN General
ChatCounselor [209] Psy DRG Vicuna Finetuning EN Mental Health
CharacterChat [210] Per, M DRG LLaMA Finetuning EN General
Fu et al. [211] E CE, DRG ChatGPT, T5 FS, Finetuning EN General
†S–Semantics, Per–Persona, Psy–Psychology, E–Emotion, M–Memory
* DS–Dialogue Summarization, NLU–Natural Language Understanding, DST–Dialogue State Tracking, PD–Proactive Dialogues, DDA–Dialogue Discourse
Analysis, DE–Dialogue Evaluation, DRG–Dialogue Response Generation, DQA–Dialogue Question Answering, DU–Dialogue Understanding, CE–Causality
Explanation
@@
One crucial aspect is equipping dialogue systems with human-like perception abilities and reasoning abilities. These abilities play a key role in comprehending the internal status of the dialogue context, which stands for any information that can be reasoned and deduced from semantic context during the conversation, benefiting several downstream dialogue tasks: dialogue understanding [206], dialogue state tracking [199, 198, 187], emotional and personalized dialogue generation [5, 208, 211], dialogue summarization [201] and so on. Regarding various prompting methods to reason different aspects of internal status exhibited in the dialogue context, we divide these methods into two categories: 1) semantic perception, which directly extracts useful information from the dialogue context in the required format; 2) semantic reasoning, which reason and deduce some underlying information beneath the dialogue context. Table V shows recent progress in this field.
5.1.1 Semantic Perception
@@
There are some works that attempt to evaluate the performance of LLM at various dialogue understanding tasks such as NLU, DST, and dialogue discourse analysis tasks. For example, Pan et al. [197] and Michael et al. [198] evaluate the capability of ChatGPT to compress the dialogue context into the semantic frame (structured output of NLU) and belief states (structured output of DST) in the context of TOD, respectively. Furthermore, Fan et al. [202] investigate ChatGPT’s performance in three discourse analysis tasks: topic segmentation, discourse relation recognition, and discourse parsing, of varying difficulty levels. The key is to compress the dialogue context, the unstructured text into structured data, to meet the requirement of downstream applications or tasks. On the other hand, another line of work tries to compress the lengthy dialogue context while keeping important semantic information, particularly in dialogue summarization tasks [201].
5.1.2 Semantic Reasoning
@@
Besides semantics information, it is widely acknowledged that dialogue contexts generally convey a lot of information about the user status from a linguistic perspective [212, 213]. Specifically, the linguistic cues underlying dialogue context have been shown to be an effective means of revealing the emotions [214], personality traits [212], psychological characteristics [213], and other relevant information of users [215]. Consequently, recognizing and understanding these cues exhibited in the context of dialogues becomes crucial to comprehend user intentions and status [5]. By doing so, a dialogue system can generate responses that align with the user’s expectations, meet their unique needs and preferences, and ultimately strive towards constructing a human-like conversational experience [5, 7].
Each linguistic cue has garnered considerable scholarly attention, resulting in a wealth of research and a variety of methodological approaches inspired by them. Specifically, Psychology: Liu et al. [209] propose ChatCounselor to provide mental counseling while Qiu et al. [207] utilize ChatGPT to rewrite the QA dataset into multi-turn mental dialogues. Emotion: Fu et al. [211] integrates commonsense-based causality explanation, considering both the user’s perspective (user’s desires and reactions) and the system’s perspective (system’s intentions and reactions). Zhao et al. [206] evaluate the performance of ChatGPT on emotional dialogue in terms of various understanding and generation tasks such as emotion recognition, emotion cause recognition, dialogue act classification (emotion dialogue understanding), empathetic response generation and emotion support generation. In order to navigate limited data and non-standardized training in emotional dialogues, the ExTensible Emotional Support dialogue dataset (ExTES) is curated by utilizing the in-context learning potential of ChatGPT that synthesizes human insights. Persona: Tu et al. [210] present CharacterChat, an extensive conversational system designed to provide social support. CharacterChat is comprised of a conversational model that is guided by personas and memory, alongside an interpersonal matching plugin model. This plugin model is responsible for identifying and assigning the most suitable supporters from the MBTI-1024 Bank to individuals based on their specific personas. All-in-one: Instead of considering different cues in isolation, Wang et al. [5] seamlessly blend them all into one cohesive conversational flow, regarding different aspects of cues as part of user status and prompt the LLMs to reason user status exhibited in the dialogue context, aiming to generate more helpful and acceptable responses for users.
@@
Besides different linguistic cues, it is noteworthy that users may occasionally fail to furnish adequate semantic information within the dialogue context, resulting in ambiguous queries or unreasonable requests. Both of these are considered key aspects of a conversational agent’s proactivity [216, 196]. To trigger the proactivity of LLMs, ProCoT [196] is proposed to augment LLMs with the goal planning capability over descriptive reasoning chains, specifically focusing on three key aspects of proactive dialogues: clarification, target-guided, and non-collaborative dialogues.
5.2 External Sources
@@
To ensure the delivery of up-to-date, informative, trustworthy, and personalized responses, it is necessary to learn and plan how to interact with external different knowledge sources [217]. In this context, the integration of these knowledge sources not only enhances the conversational prowess of the LLM-based dialogue system but also enables them to remain adaptable and aligned with the dynamic nature of the information they handle, ultimately elevating their utility and reliability in real-world scenarios. In this section, we mainly present how LLM-based dialogue systems integrate with these knowledge sources and plan their usage.
  • [leftmargin=*,topsep=0pt,itemsep=0pt,parsep=0pt]
@@
  • Database. TOD mostly relies on one specific database such as the restaurant or flight database to answer the user query, providing the up-to-date and necessary information to accomplish the task [199, 187]. Recently, Hudevceke et al. [199] and Zhang et al. [187] consecutively evaluate the performance of ChatGPT on the E2E TOD. In detail, Hudevceke et al. [199] introduce a novel approach for dialogue state tracking, where belief states are derived from dialog history by utilizing slot descriptions as prompts, subsequently leading to database entry retrieval and response generation. In contrast, SGP-TOD [187] employs slot names and value examples, rather than slot descriptions, as prompts to facilitate frozen LLMs in generating belief states, thereby reducing human effort. In addition, they offer a policy skeleton within the Policy Prompter to guide LLMs in producing responses that comply with the predefined dialog policy, resulting in better performance than IT-LLM [199]. In order to maximize the utilization of databases within dialogue systems, ChatDB [218] employs large-scale models to generate SQL instructions during the dialogue process.
@@
  • Memory. The importance of memory knowledge sources in the context of large language models (LLMs) for open-domain dialogue systems is paramount. These memory sources enable LLMs to retain and access relevant information from previous interactions, facilitating more coherent and contextually aware responses in long-term conversations. Effective memory mechanisms are crucial for a wide range of applications, including personal companion systems, psychological counseling, and secretarial assistance, where sustained interactions and the ability to recall past information are vital for providing a rich and personalized user experience [219, 220, 221]. Particularly, Zhong et al. [220] introduce MemoryBank, a novel memory mechanism inspired by the Ebbinghaus Forgetting Curve theory. MemoryBank allows LLMs to summon and selectively update memories, adapting to a user’s personality over time. It demonstrates the utility of MemoryBank in a long-term AI companion scenario. Wang et al. [221] introduce a recursive memory generation approach, where LLMs are trained to memorize small dialogue contexts and generate new memory as conversations progress, aiming to improve consistency in long-context conversations. Furthermore, Lu et al. [222] present MemoChat, a pipeline for refining instructions to enable LLMs to use self-composed memos to maintain consistency in long-range conversations. It focuses on the iterative “memorization-retrieval-response” process and incorporates tailored instructions. These works collectively contribute to addressing the challenge of long-term memory in LLMs with distinct approaches and innovations.
@@
  • Internet. One key notorious problem of LLMs is the hallucinations [6] [^10] Most works about the hallucination can be found in § 6.2.. Prior studies demonstrate that external document sources can be used to validate the factuality of generated response [223] and thus reduce the hallucinations [171, 6]. In addition, document sources can be used to store up-to-date information, enabling the LLM to sync the latest world knowledge such as New Bing. While responses of LLM-based DS are more likely to include information not present in the relevant document, possibly including the presence of hallucinations, they are preferred by human judges compared with other baselines without using LLM and even human responses [224]. To leverage enhancing the quality and reliability of responses together, WebGLM [225] extends LLMs with web search and retrieval capabilities, demonstrating the efficiency and effectiveness in question-answering tasks, while Semnani et al. [226] mainly leverage documents from Wikipedia to enhance the factuality of LLM-based DS, focusing on providing accurate and verified information with additional fact-checking and refinement steps. Furthermore, there are some works that explore the relationship between multiple documents [227] or queries [228] to solve one complex question requiring multi-hop reasoning.
@@
  • Profile. Salemi et al. [7] first introduce the LaMP benchmark — a novel benchmark for training and evaluating LLMs for producing personalized outputs. LaMP offers a comprehensive evaluation framework with diverse language tasks and multiple entries for each user profile, consisting of three personalized understanding and four personalized generation tasks.
@@
  • Others. Besides these common knowledge sources, there are other knowledge sources in a different organization or structure that also play a key role during response generation such as API [229, 230, 231], and Knowledge Graph [205].
@@
  • All-in-one. Most existing LLM-based DS either focus on a single knowledge source or overlook the relationship between multiple sources of knowledge, which may result in generating inconsistent or even paradoxical responses [232, 233]. To incorporate multiple sources of knowledge and relationships between them, Wang et al. [217] propose SAFARI, a novel framework that takes advantage of the exceptional capabilities of large language models (LLM) in planning, understanding, and incorporating in both supervised and unsupervised settings. Specifically, SAFARI decouples the knowledge grounding into multiple knowledge sources selection and response generation, which allows easy extension to various knowledge sources including the possibility of not using any sources.
5.3 Internal Reasoning with External Sources

### v2

@@ 4.3. Fusion between LLM and DM
To conclude, several trends can be drawn from the fusion results at Table 5. First of all, there is a discernible temporal trend towards an increased prevalence of open-sourced LLMs, particularly within the domain of dialogue models. Secondly, this trend is accompanied by continuous growth in both training data and parameter configurations, although most released models tend to be around 7B or 13B parameters due to the high computational costs associated with inference. In addition, there is a noteworthy extension in the window length, denoting an enhanced ability to process and generate longer sequences of text. To further improve the utility and safety of dialogue models, extensive instruction fine-tuning using conversational data is increasingly integrated, and their multilingual capabilities continue to evolve.
Table 7. Performance Comparison of Different Dialogue Methods across TOD and OOD Tasks. Note: Blue-highlighted rows indicate Unified Dialogue Systems; yellow-highlighted rows indicate End-to-End Dialogue Models.
Task Type Dataset Method Type Method Metric Performance
ToD-NLU MSDialog (Qu and et al., 2018) Tra. DS BERT-base (Devlin and et al., 2019) P, F1-Micro, F1-Macro 81.0, 83.1, 74.3
T5-base 83.9, 86.8, 77.1
LLM-Prompting GPT-4o 54.6, 95.8, 42.4
LLM-Finetuning SOLID (Askari and et al., 2024) 84.2, 87.9, 78.6
ToD-DST Multi-Woz 2.1 (Eric and et al., 2020) Tra. DS SimpleTOD (Hosseini-Asl et al., 2020) JGA 54.15
TOATOD (Bang and et al., 2023a) 54.79
DSGFNet (Feng and et al., 2022) 56.79
LLM-Prompting GPT-4o 62.59
Llama2-13B 49.28
VICUNA-13B-V1.5 55.41
LLM-Finetuning FNCTOD (Li and et al., 2024) 59.54
SERIDST (Lee and et al., 2024) 56.96
GPT-3.5 61.31
ToD-DPL Multi-Woz 2.0 (Budzianowski and et al., 2018) Tra. DS GALAXY (He and et al., 2022a) Inform, Comb 93.5, 110.4
DiactTOD (Wu and et al., 2023a) 94.6, 97.1
Unified DS UniDS (Zhao et al., 2022) 90.3, 104.12
LLM-Prompting GPT-4 85.2, 79.1
Llama2 70B 54.3, 38.9
Llama2 13B 37.1, 33.6
LLM-Finetuning AutoTOD (Xu and et al., 2024a) 85.2, 79.1
GPT-3.5 62.5, 52.7
ToD-NLG Multi-Woz 2.1 Tra. DS UBAR (Yang et al., 2021) SUCCESS, BLEU 89.1, 14.8
GALAXY (He and et al., 2022a) 84.9, 20.8
Unified DS HyKnow (Gao and et al., 2021) 69.4, 18.0
LLM-Prompting GPT-3.5 82.8, 9.3
GPT-4 84.4, 10.4
Llama2-70B 69.8, 7.8
LLM-Finetuning AutoTOD (Xu and et al., 2024a) 84.9, 7.8
ToD-E2E Multi-Woz 2.0 Tra. DS PPTOD (Su and et al., 2021) BLEU, Inform, Success 18.1, 62.9, 79.4
SPACE-3 (He and et al., 2022b) 19.3, 95.3, 88.0
Unified DS UniDS (Zhao et al., 2022) 18.72, 80.3, 80.05
LLM-Prompting GPT-3.5 19.4, 87.0, 79.1
GPT-4 10.4, 91.7, 84.4
Llama2-70B 7.8, 73.3, 69.9
LLM-Finetuning InstructTODS (Chung and et al., 2023) 9.34, 90.7, 76.2
OOD ESConv (Liu and et al., 2021) Tra. DS KEMI (Deng and et al., 2023b) BL-2, RG-L, Dist-3 2.7, 2.9, 2.3
TransESC (Zhao and et al., 2023) 2.6, 2.8, 2.5
LLM-Prompting Llama-3-70B 2.1, 3.0, 2.6
GPT-3.5-turbo 3.4, 2.5, 3.5
LLM-Finetuning COOPER (Cheng and et al., 2024) 2.9, 5.4, 3.4
The fusion of LLMs and DMs offers transformative benefits, particularly in enhancing performance and improving generalization. Table 7 presents the performance of different types of dialogue system on some representative benchmarks. There are several trends can be drawn from the results. First of all, Traditional task-specific methods (Tra. DS) generally perform well in specific tasks, but they often lack flexibility and adaptability across different dialogue challenges. Secondly, recent LLMs, such as GPT-3.5 and GPT-4, achieve exceptional performance and, in some cases, even outperform traditional task-specific methods (Tra. DS) in a zero-shot setting, in both TOD or ODD tasks like ToD-DST, ToD-E2E and ESConv. Thirdly, some finetuned small sized LLMs can outperform relatively larger sized LLMs on certain tasks, such as AutoToD outperforms GPT-3.5 in ToD-DPL. Overall, by leveraging extensive pretraining, LLMs develop vast world knowledge, nuanced reasoning abilities, and robust generalization across diverse natural language tasks. This comprehensive pretraining significantly reduces the need for task-specific fine-tuning — not only in dialogue systems but across nearly all language-based applications — enabling these models to function more like human assistants rather than solely dialogue systems. However, this kind of fusion is not without limitations. The overall performance of an integrated dialogue model is heavily contingent upon the underlying LLM; thus, any biases or limitations inherent in the base model can be transferred to the dialogue system. Moreover, language dominance exacerbates existing inequities: models for widely represented languages like English and Chinese benefit from abundant training data, whereas underrepresented languages (e.g., Sherkala) struggle with sparse data and limited fine-tuning efforts. This imbalance risks cultural homogenization, as dominant languages increasingly shape dialogue norms and priorities. Ultimately, while the fusion of LLMs and DMs marks a significant leap forward, addressing these challenges is essential to ensure that dialogue systems are both equitable and culturally diverse.
4.4. Summary
Over the decades, dialogue systems have evolved from rule-based and modular architectures — where separate components handled tasks like natural language understanding, dialogue management, and response generation — to unified end-to-end systems that leverage the power of deep learning and large-scale pretraining. The three key waves of fusions, starting from fusions within TOD, then fusion between TOD and ODD, and finally LLM with the DM, exactly reflect the fusion from bottom to top, being more and more scalable and generaliable across data, tasks and models. One of key lessons from the history of dialogue system is exactly the Bitter Lesson [^11] http://www.incompleteideas.net/IncIdeas/BitterLesson.html, which progress has been driven by scaling computation and data, not human-centric engineering, less structure, more intelligence. These difference levels of fusions can be regarded as a progress keeping reduce human-centric engineering but leverage scaling computation and data.
Looking forward, the future of dialogue systems will likely continue to be driven by the bitter lesson’s core tenet: computation and learning, not human ingenuity, are the engines of intelligence. As models become increasingly unified — blurring the lines between different tasks, languages and modalities – they promise more natural, context-aware, and adaptable interactions. However, these new fusions also bring challenges, including the new learning paradigm, unified data format and even new model architecture as we observed in the development of dialogue system. Ultimately, by embracing the lessons of the past and leveraging the immense power of computation, the next generation of dialogue systems can become more robust, versatile, and human-like in their understanding and interactions.
5. LLM-based Dialogue System
Nowadays, it is not a formidable challenge for LLM-based dialogue systems to pass the esteemed Turing Test (Jones and et al., 2025). More importantly, as dialogue systems and backbone language models evolve, the boundaries between tasks, models, and even languages are likely to blur. A fundamental capability of such models will be their ability to mimic human cognitive processes and interact effectively with external environments, in order to get more helpful, harmless and natural human-like response. This raises the critical question of how models can employ cognitive reasoning internally and simultaneously interact with environment externally. The unification of these two abilities will be very foundation to creating models that not only reason current state but also act as adaptive problem-solvers in complex real-world contexts.
Figure 3 shows the three major conversational flows for the LLM-based dialogue system. First of all, as shown in left side, the LLM-based dialogue system need to understand the user intention and interprets subtle cues or implicit expectations with greater nuance by leveraging exceptional reasoning capabilities of LLMs (§ 5.1), similar with the recently released GPT 4.5121212https://openai.com/index/introducing-gpt-4-5/. Secondly, it should interact with various external knowledge sources — such as long-term memory systems, private databases, and domain-specific services — to generate more informed and accurate responses with reduced hallucinations and up-to-date information (§ 5.2). Lastly, the system carries out these operations through a multi-step process, enabling more sophisticated decision-making and problem-solving (§ 5.3). Together, these three critical capabilities have become increasingly attractive and crucial for empowering dialogue systems, which can be utilized in complementary to suit the diverse and complicated needs in practice.
@@ 5. LLM-based Dialogue System
Figure 3. A novel unified perspective of LLM-based Dialogue System. There are three conversational flows: 1) response based on internal reasoning and dialogue context; 2) response based on external acting and dialogue context; and 3) response based on internal reasoning, external acting, and dialogue context.
5.1. Internal Reasoning
@@
Table 8. Summary of research works exploring LLMs for various dialogue tasks.
Method(s) Cue(s)† Type Task(s)* LLM(s) Settings Language Domain Pan et al. (Pan and et al., 2023) S TOD NLU ChatGPT ZS EN General Michael et al. (Heck and et al., 2023) S DST ChatGPT ZS EN General IT-LLM (Hudeček and Dušek, 2023) S DST, NLG ChatGPT ZS, FS EN General SGP-TOD (Zhang and et al., 2023c) S DST, DPL, NLG ChatGPT ZS EN General GDP-ZERO (Yu and et al., 2023) S DPL ChatGPT FS EN PersuasionForGood Qin et al. (Qin et al., 2023) S ODD DS ChatGPT ZS EN General ProCoT (Deng and et al., 2023c) S PD ChatGPT, Vicuna ZS, FS EN General Fan et al. (Fan and et al., 2023) S DDA ChatGPT ZS, FS CN, EN General Huynh et al. (Huynh and et al., 2023) S DE GPT3, BLOOM FS EN General Mendonça et al. (Mendonça and et al., 2023) S DE ChatGPT ZS Multilingual General Bang et al. (Bang and et al., 2023b) S TOD, ODD DRG ChatGPT ZS EN General CueCoT (Wang and et al., 2023b) Per, Psy, E ODD DQA, DRG ChatGPT, ChatGLM ZS, FS CN, EN General Zhao et al. (Zhao et al., 2023) E DU, DRG ChatGPT ZS EN General SMILE (Qiu and et al., 2023) Psy ODD DRG CN Mental Health ExTES (Zheng et al., 2023) E ODD DRG LLaMA Finetuning EN General ChatCounselor (Liu et al., 2023b) Psy DRG Vicuna Finetuning EN Mental Health CharacterChat (Tu and et al., 2023) Per, M DRG LLaMA Finetuning EN General Fu et al. (Fu and et al., 2023) E CE, DRG ChatGPT, T5 FS, Finetuning EN General †S–Semantics, Per–Persona, Psy–Psychology, E–Emotion, M–Memory * DS–Dialogue Summarization, NLU–Natural Language Understanding, DST–Dialogue State Tracking, PD–Proactive Dialogues, DDA–Dialogue Discourse Analysis, DE–Dialogue Evaluation, DRG–Dialogue Response Generation, DQA–Dialogue Question Answering, DU–Dialogue Understanding, CE–Causality Explanation
@@
One crucial aspect is equipping dialogue systems with human-like perception abilities and reasoning abilities (Wang et al., 2025a). These abilities play a key role in comprehending the internal status of the dialogue context, which stands for any information that can be reasoned and deduced from semantic context during the conversation, benefiting several downstream dialogue tasks: dialogue understanding (Zhao et al., 2023; Zhang et al., 2024), dialogue state tracking (Hudeček and Dušek, 2023; Heck and et al., 2023; Zhang and et al., 2023c, 2025), emotional and personalized dialogue generation (Wang and et al., 2023b; Zheng et al., 2023; Fu and et al., 2023), dialogue summarization (Qin et al., 2023) and so on. Regarding various prompting methods to reason different aspects of internal status exhibited in the dialogue context, we divide these methods into two categories: 1) semantic perception, which directly extracts useful information from the dialogue context in the required format; 2) semantic reasoning, which reason and deduce some underlying information beneath the dialogue context to better understand user intention. Table 8 shows recent progress in this field.
5.1.1. Semantic Perception
@@
There are some works that attempt to evaluate the performance of LLM at various dialogue understanding tasks such as NLU, DST, and dialogue discourse analysis tasks. For example, Pan et al. (Pan and et al., 2023) and Michael et al. (Heck and et al., 2023) evaluate the capability of ChatGPT to compress the dialogue context into the semantic frame (structured output of NLU) and belief states (structured output of DST) in the context of TOD, respectively. Furthermore, Fan et al. (Fan and et al., 2023) investigate ChatGPT’s performance in three discourse analysis tasks: topic segmentation, discourse relation recognition, and discourse parsing, of varying difficulty levels. The key is to compress the dialogue context, the unstructured text into structured data, to meet the requirement of downstream applications or tasks. On the other hand, another line of work tries to compress the lengthy dialogue context while keeping important semantic information, particularly in dialogue summarization tasks (Qin et al., 2023).
5.1.2. Semantic Reasoning
@@
Besides semantics information, it is widely acknowledged that dialogue contexts generally convey a lot of information about the user status from a linguistic perspective (Tausczik and Pennebaker, 2010). Specifically, the linguistic cues underlying dialogue context have been shown to be an effective means of revealing the emotions (Ekman, 1971), personality traits, psychological characteristics (Tausczik and Pennebaker, 2010), and other relevant information of users (Turney, 2002). Consequently, recognizing and understanding these cues exhibited in the context of dialogues becomes crucial to comprehend user intentions and status (Wang and et al., 2023b). By doing so, a dialogue system can generate responses that align with the user’s expectations, meet their unique needs and preferences, and ultimately strive toward constructing a human-like conversational experience (Wang and et al., 2023b; Salemi et al., 2023). Each linguistic cue has garnered considerable scholarly attention, resulting in a wealth of research and a variety of methodological approaches inspired by them. Specifically, there are three major types of cues center around the user and there are some studies attempt to seamlessly integrates these multiple linguistic cues into a cohesive conversational framework.
  • Psychology. In the realm of mental health and counseling, it is crucial to consider the psychological state of user and then tailor the response to provide mental support (Sun and et al., 2021). Recently, Liu et al. (2023) introduced ChatCounselor, a model designed to provide mental health counseling. Meanwhile, Qiu et al. (2023) leveraged ChatGPT to transform a QA dataset into multi-turn mental health dialogues, enabling more effective fine-tuning on specialized datasets.
  • Emotion. Emotional intelligence has always been a crucial feature of dialogue systems, enabling them to understand and generate more empathetic and contextually appropriate responses (Wang et al., 2023b). For examples, Fu et al. (Fu and et al., 2023) integrates commonsense-based causality explanation, considering both the user’s perspective (user’s desires and reactions) and the system’s perspective (system’s intentions and reactions). Zhao et al. (Zhao et al., 2023) evaluate the performance of ChatGPT on emotional dialogue in terms of various understanding and generation tasks such as emotion recognition, emotion cause recognition, dialogue act classification (emotion dialogue understanding), empathetic response generation and emotion support generation. In order to navigate limited data and non-standardized training in emotional dialogues, the ExTensible Emotional Support dialogue dataset (ExTES) is curated by utilizing the in-context learning potential of ChatGPT that synthesizes human insights.
  • Personality. The way one speaks reflects their personality. It is straightforward to infer the user persona from textual information especially the dialogue context (Wang and et al., 2023b). In detail, Tu et al. (Tu and et al., 2023) present CharacterChat, an extensive conversational system designed to provide social support. CharacterChat is comprised of a conversational model that is guided by personas and memory, alongside an interpersonal matching plugin model. This plugin model is responsible for identifying and assigning the most suitable supporters from the MBTI-1024 Bank to individuals based on their specific personas.
  • Unification. Instead of considering different cues in isolation, Wang et al. (Wang and et al., 2023b) seamlessly blend them all into one cohesive conversational flow, regarding different aspects of cues as part of user status and prompt the LLMs to reason user status exhibited in the dialogue context, aiming to generate more helpful and acceptable responses for users.
@@
Besides different linguistic cues, it is noteworthy that users may occasionally fail to furnish adequate semantic information within the dialogue context, resulting in ambiguous queries or unreasonable requests. Both of these are considered key aspects of a conversational agent’s proactivity (Deng et al., 2023; Deng and et al., 2023c). To trigger the proactivity of LLMs, ProCoT (Deng and et al., 2023c) is proposed to augment LLMs with the goal planning capability over descriptive reasoning chains, specifically focusing on three key aspects of proactive dialogues: clarification, target-guided, and non-collaborative dialogues.
5.2. External Acting
@@
To ensure the delivery of up-to-date, informative, trustworthy, and personalized responses, it is necessary to learn and plan how to interact with external different knowledge sources (Wang and et al., 2023c; Wang et al., 2025d). Therefore, the integration of these knowledge sources not only enhances the conversational prowess of the LLM-based dialogue system but also enables them to remain adaptable and aligned with the dynamic nature of the information they handle, ultimately elevating their utility and reliability in real-world scenarios. In this section, we mainly present how LLM-based dialogue systems interact with external environment to integrate with different knowledge sources and plan their usage.
@@
  • Database. TOD mostly relies on one specific database such as the restaurant or flight database to answer the user query, providing the up-to-date and necessary information to accomplish the task (Hudeček and Dušek, 2023; Zhang and et al., 2023c). Recently, Hudevceke et al. (Hudeček and Dušek, 2023) and Zhang et al. (Zhang and et al., 2023c) consecutively evaluate the performance of ChatGPT on the E2E TOD. In detail, Hudevceke et al. (Hudeček and Dušek, 2023) introduce a novel approach for dialogue state tracking, where belief states are derived from dialog history by utilizing slot descriptions as prompts, subsequently leading to database entry retrieval and response generation. In contrast, SGP-TOD (Zhang and et al., 2023c) employs slot names and value examples, rather than slot descriptions, as prompts to facilitate frozen LLMs in generating belief states, thereby reducing human effort. In addition, they offer a policy skeleton within the Policy Prompter to guide LLMs in producing responses that comply with the predefined dialog policy, resulting in better performance than IT-LLM (Hudeček and Dušek, 2023). In order to maximize the utilization of databases within dialogue systems, ChatDB (Hu et al., 2023) employs large-scale models to generate SQL instructions during the dialogue process. Moreover, recent studies try to formulate the API call as tool utilization problems such as AppBench (Wang et al., 2024) and DilogTool (Wang et al., 2025b).
@@
  • Memory. The importance of memory knowledge sources in the context of large language models (LLMs) for open-domain dialogue systems is paramount. These memory sources enable LLMs to retain and access relevant information from previous interactions, facilitating more coherent and contextually aware responses in long-term conversations. Effective memory mechanisms are crucial for a wide range of applications, including personal companion systems, psychological counseling, and secretarial assistance, where sustained interactions and the ability to recall past information are vital for providing a rich and personalized user experience (Lee and et al., 2023; Zhong and et al., 2023; Wang and et al., 2023e). Particularly, Zhong et al. (Zhong and et al., 2023) introduce MemoryBank, a novel memory mechanism inspired by the Ebbinghaus Forgetting Curve theory. MemoryBank allows LLMs to summon and selectively update memories, adapting to a user’s personality over time. It demonstrates the utility of MemoryBank in a long-term AI companion scenario. Wang et al. (Wang and et al., 2023e) introduce a recursive memory generation approach, where LLMs are trained to memorize small dialogue contexts and generate new memory as conversations progress, aiming to improve consistency in long-context conversations. Furthermore, Lu et al. (Lu and et al., 2023) present MemoChat, a pipeline for refining instructions to enable LLMs to use self-composed memos to maintain consistency in long-range conversations. It focuses on the iterative “memorization-retrieval-response” process and incorporates tailored instructions. These works collectively contribute to addressing the challenge of long-term memory in LLMs with distinct approaches and innovations.
@@
  • Internet. One key notorious problem of LLMs is the hallucinations (Huang et al., 2021) [^13] Most works about the hallucination can be found in § 6.4.. Prior studies demonstrate that external document sources can be used to validate the factuality of generated response (Wang and et al., 2023a) and thus reduce the hallucinations (Shuster et al., 2021; Huang et al., 2021). In addition, document sources can be used to store up-to-date information, enabling the LLM to sync the latest world knowledge such as New Bing. While responses of LLM-based DS are more likely to include information not present in the relevant document, possibly including the presence of hallucinations, they are preferred by human judges compared with other baselines without using LLM and even human responses (Braunschweiler and et al., 2023). To leverage enhancing the quality and reliability of responses together, WebGLM (Liu and et al., 2023b) extends LLMs with web search and retrieval capabilities, demonstrating the efficiency and effectiveness in question-answering tasks, while Semnani et al. (Semnani and et al., 2023) mainly leverage documents from Wikipedia to enhance the factuality of LLM-based DS, focusing on providing accurate and verified information with additional fact-checking and refinement steps. Furthermore, there are some works that explore the relationship between multiple documents (Trivedi and et al., 2023) or queries (Wang and et al., 2025) to solve one complex question requiring multi-hop reasoning.
@@
  • Profile. Salemi et al. (Salemi et al., 2023) first introduce the LaMP benchmark — a novel benchmark for training and evaluating LLMs for producing personalized outputs. LaMP offers a comprehensive evaluation framework with diverse language tasks and multiple entries for each user profile, consisting of three personalized understanding and four personalized generation tasks.
@@
  • Others. Besides these common knowledge sources, there are other knowledge sources in a different organization or structure that also play a key role during response generation such as API (Li et al., 2023; Patil et al., 2023; Liang and et al., 2023), and Knowledge Graph (Bang and et al., 2023b).
@@
  • Unification. Most existing LLM-based DS either focus on a single knowledge source or overlook the relationship between multiple sources of knowledge, which may result in generating inconsistent or even paradoxical responses (Bodhisattwa and et al., 2020; Fu and et al., 2022). To incorporate multiple sources of knowledge and relationships between them, Wang et al. (Wang and et al., 2023c) propose SAFARI, a novel framework that takes advantage of the exceptional capabilities of large language models (LLM) in planning, understanding, and incorporating in both supervised and unsupervised settings. Specifically, SAFARI decouples the knowledge grounding into multiple knowledge sources selection and response generation, which allows easy extension to various knowledge sources including the possibility of not using any sources.
5.3. Internal Reasoning with External Acting